Brighton Services repairs locomotive engines. It employs 100 full-time workers at $14 per hour. Despite operating at capacity, last year's performance was a great disappointment to the managers. In total, 10 jobs were accepted and completed, incurring the following total costs.

Direct materials $ 1,045,400
Direct labor 4,200,000
Manufacturing overhead 975,000

Of the $975,000 manufacturing overhead, 40 percent was variable overhead and 60 percent was fixed.

This year, Brighton Services expects to operate at the same activity level as last year, and overhead costs and the wage rate are not expected to change. For the first quarter of this year, Brighton Services completed two jobs and was beginning the third (Job 103). The costs incurred follow.

Job Direct Materials Direct Labor
101 $ 138,200 $ 490,000
102 103,000 313,000
103 95,000 195,200
Total manufacturing overhead 272,200
Total marketing and administrative costs 120,000

You are a consultant associated with Lodi Consultants, which Brighton Services has asked for help. Lodi's senior partner has examined Brighton Services's accounts and has decided to divide actual factory overhead by job into fixed and variable portions as follows.

Actual Manufacturing Overhead
Variable Fixed
101 $ 30,900 $ 105,000
102 28,500 89,200
103 5,600 13,000
$ 65,000 $ 207,200

In the first quarter of this year, 40 percent of marketing and administrative cost was variable and 60 percent was fixed. You are told that Jobs 101 and 102 were sold for $855,000 and $570,000, respectively. All over- or underapplied overhead for the quarter is written off to Cost of Goods Sold.

Required:

a. Present in T-accounts the actual manufacturing cost flows for the three jobs in the first quarter of this year.

b. Using last year's overhead costs and direct labor-hours as this year's estimate, calculate predetermined overhead rates per direct labor-hour for variable and fixed overhead.

c. Present in T-accounts the normal manufacturing cost flows for the three jobs in the first quarter of this year. Use the overhead rates derived in requirement (b).

d. Calculate operating profit (loss) for the first quarter of this year under actual and normal costing systems.
Answer #1

rA)   Material inventory                                                  

beg bal00

0138,200101 dm

0103,000102 dm

095,000103 dm

00
end bal
336,200
wages payable

beg bal00


0490,000101 dl

0313,000102dl

0195,200103 DL

00
end bal
998,200
variable manufacturing overhead

actual65,00030,900101 variable

028,500102 v

05,600103 v

00
end bal


 fixed MOH

  

actual207,200105,000101 fixed

089,200102 fixed

013,000103 fixed

00
end bal


WIP inventory

beg bal00
total dm336,200764,100total 101 finished goods
total DL998,200533,700total 102 ''  ''
total variable overhead65,0000
total fixed overhead207,2000

00




101  138,200+ 490,000+ 105,000+ 30,900= 764,100 

102   103,000+ 313,000+ 89,200+  28,500= 533,700

from the two chats above


finished goods inventory

beg bal00
total 101 finished goods764,1001,297,800Cost of goods sold
total 102 '' ''533,7000

00
end bal


764,100+533,700=  1,297,800


COGS (cost of goods sold)


beg bal00
finished goods1,297,8000

00
end bal1,297,800
Answer #2

rB) 975,000* 40%= 390,000 variable

      975,000*60%= 585,000  fixed

4,200,000/14= 300,000 dl hours


variable 390,000/300,000=1.3

fixed 585,000/300,000= 1.95
Answer #3

rC) Material inventory

beg bal00

0138,200101 dm

0103,000102 DM

095,000103 DM

00
end bal
336,200

wages payable

beg bal00

0490,000101 DL

0313,000102 DL

0195,200103 DL

00
end bal
998,200

variable manufacturing overhead (MOH)

actual65,00045,500101 variable
overapplied27,69029,064

102 variable


018,126103 variable

00
end bal


101 v: 490,000/14 * 1.3=  45,500

102 v: 313,000/14 * 1.3=  29,064.285714286 or 29,064

103 v:  195,200/14 * 1.3= 18,125.714285714 or 18,126 round 1 since above 5

total variable 45,500+ 29,064+ 18,126= 92,690

overapplied 92,690-65,000=  27,690


Fixed MOH

actual207,20068,250101 fixed

043,596102 fixed

027,189103 fixed

068,165underapplied

00
end bal


101 f: 490,000/14 *1.95= 68,250

102 f: 313,000/14 *1.95= 43,596.428571429 or 43,596

103f: 195,200/14 *1.95= 27,188.571428571 or 27,189

total fixed 68,250+ 43,596+ 27,189= 139,035

underapplied 207,200-139,035= 68,165


WIP inventory

beg bal00
total DM336,200741,950

total 101 

finished goods

total DL998,200488,660

total 102 

finished goods

total fixed MOH139,0350

total variable 

MOH

92,6900

0
0
end bal


101 138,200+ 490,000+ 45,500(variable rC)+ 68,250(fixed rC)= 741,950

102 103,000+ 313,000+ 29,064+ 43,596= 488,660


Finished goods inventory

beg bal00

total 101 

finished goods

741,9501,230,610COGS

total 102

finished goods

488,6600

00
end bal


741,950+ 488,660= 1,230,610


COGS

beg bal00

finished

goods

1,230,6100

00
end bal


under or overapplied overhead

beg bal00
underapplied68,16527,690overapplied

00
end bal40,475
0

68,165-27,690= 40,475
Answer #4

rD) Actual

sales revenue 855,000+570,000= 1,425,000

less COGS: 1,297,800 from rA )COGS 

gross margin: 1,425,000- 1,297,800= 127,200

Less:(under-) overapplied overhead: 0 no t-chart for part A

Marketing and administrative cost: 120,000 from problem explination

Operating profit (loss): 127,200- 120,000= 7,200


Normal

sales revenue 855,000+570,000= 1,425,000

less COGS:  1,230,610   from rC )COGS 

gross margin: 1,425,000- 1,230,610= 194,390

Less:(under-) overapplied overhead: 40,475 from part C t0chart

Marketing and administrative cost:: 120,000

Operating profit (loss): 194,390- 40,475- 120,000= 33,915
Answer #5

A) Present in T-accounts the actual manufacturing cost flows for the three jobs in the first quarter of this year.

MATERIALS INVENTORY
AMT $ AMT($)
Beginning balance 0 Direct Materials (101) 138200
0 Direct Materials (102) 103000
0 Direct Materials(103) 95000
Closing Balance 0 336200
Wages payable
AMOUNT $ AMOUNT $
Beginning balance 0 Direct Labour
101 490000
102 313000
103 195200
Closing Balance 0

998200

VARIABLE MANUFACTURING OVERHEAD
AMT $ AMT $
Beginning Balance 0 101 30900
102 28500
103 5600
Closing Balance 65000

65000

FIXED MANUFACTURING OVERHEAD
AMT $ AMT $
Beginning balance 0 101 105000
102 89200
103 13000
closing balance 207200 207200
Work in progress Inventoty
Amount $ Amount $
beginning balance 0 Total finished goods
TotaL Direct materials 336200 101 764100*
TotaL Direct Labour      998200 102 533700*
total variable Overhead 65000
total fixed overhead 207200
closing Bal Closing balance 308800
1606600 1606600
Finished goods Inventory
Amount $ Amount $
Beginning balance 0 Cost of goods sold 1297800
Total finished goods
101 764100
102 533700
1297800 1297800
Cost of goods Sold
beginning balance 0
Finished goods 1297800
1297800 1297800

** cost of goods sold= direct materials+Direct labour+Manufacturing overhead(fixed +variable)

cost of goods sold of 101 =138200+490000+30900+105000=764100

102 =103000+313000+28500+89200=533700

B)Using last year's overhead costs and direct labor-hours as this year's estimate, calculate predetermined overhead rates per direct labor-hour for variable and fixed overhead.

Predetermined Overhead Rate (per Direct Labour Hour)
Variable overhead rate $ 1.30
Fixed Overhead rate $1.95
Variable overhead rate= $ 975000*40/100=390000/300000= $1.30
direct labour hours = $ $ 4200000/14=300000 hours
Fixed overhead rate= 975000*60/100=585000/300000=$1.95

c)